Can someone help? I just can't figure out these "linear word problems".Thank-you
Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280 lb. mixture that is 10% protein?
Any help I appreciate.

Answer by RAY100(1637)

.14 S + .07 C = .10 *280
but s + c =280, therefore let c=280-s
.14 S +.07 ( 280-S) = 28
.14 S + 19.6 -.07 S = 28
.07 S = 8.4
S=120
and c= 280-120 = 160
check
.14 (120) + .07 ( 160) = 28
16.8 + 11.2 = 28 ok